OpenStudy (anonymous):

How many numbers are between 100 and 999 inclusive?

OpenStudy (lgbasallote):

i think you're missing a few words?

OpenStudy (anonymous):

*including 100 and 999

jimthompson5910 (jim_thompson5910):

Rule: Number of numbers from x to y = x - y + 1 Note: x < y -------------------------------------------------------- ex: Question: How many numbers are from 67 to 195? 195 - 67 + 1 = 129 Answer: So there are 129 numbers between 67 and 195 (including both 67 and 195)

OpenStudy (anonymous):

amount of numbers from a to b = b-a+1 So for this case, 999-100+1 = 900 Answer: 900 numbers